Supplementary Operations Report – January 11th

Supplemental Operations Report 11Jan2022

Name of person filing report: Atila Meszaros

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: n/a

Generator: Total run time 369.3 hours. Still being manually run at night. 20 more hours for generator maintenance, including oil filter change. Parts for maintenance are in their way.

ScienceDome Dual Split: Off

Solar—Nominal, charge controllers off at night. We are turning the generator on before the SOC reaches 70%, and if expecting cloudy days the next morning, before it reaches 80%.

Solar— VDC Last daylight hours: (information obtained from

Average 51.66

Minimum 47.7

Maximum 59.4

Notes on solar: Nothing to report

Propane Reading, station tank – 35 %

Propane Reading, director tank— 40 %

Propane Reading, intern tank— 40 %

Propane Reading, generator— 30 %

Ethanol Free Gasoline – 0 gallons/ 2 last gallons spent on the Honda. Refilled due to tomorrow

Water (Outpost tank) – 350 gallons

Perseverance rover used: No

Hours: 226.0

Beginning charge: 100

Ending charge: 100

Currently charging: Yes

Sojourner rover used: No

Hours: 166.8

Beginning charge: 100

Ending charge: 100

Currently charging: Yes

Notes on rovers: The dump stabilizer bar on Percy was removed successfully and Percy is back in action since 07 Jan.

ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0

Notes on ATVs: Oil checked on both, Honda and 350.1. Oil levels nominal.

HabCar used and why, where? Yes, for Desert Varnish, to town for a sample transport to the freezer.

CrewCar used and why, where? No

Luna used and why, where? No

Business internet: Nothing to report

General notes and comments: Nothing to report.

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: Generator turned on every evening around 7-8 PM at a 70-80% SOC. The new lock for the observatory has been set to the old code. It has been decided to only use it when crews are not actively at the station. Expedition Hab car 2022 tag sticker was placed. Observatory parts, relay and control box have been taken out and shipped to Ed for repair. Observatory dome wouldn’t close after operation, it had to be left open overnight, after a helpless attempt, the dome finally closed. Peter is contacting AstroHaven to provide support for future similar events. New mouse for the observatory laptop was delivered.

Summary of Outpost operations: Trailer containing rubbish has been covered and secured with tarp for winds and rain events.

Summary of health and safety issues: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site